# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ 2.3 Fluctuation-dissipation relations
The FDRs (51) and (52) constrain the effective action. The computation of the Lie derivatives of the metric and vector field with respect to $\beta^{\mu}$ is straightforward. We find
$\frac{\rho_{(1)}}{\beta\hbar}$ $=w_{1}\beta^{-1}u^{\rho}\nabla_{\rho}\beta-w_{2}\nabla_{\rho}u^{\rho}-x_{1}u^{\rho}\nabla_{\rho}(\beta\mu)$ (54)
$\frac{P_{(1)}}{\beta\hbar}$ $=w_{2}u^{\rho}\beta^{-1}\nabla_{\rho}\beta-w_{4}\nabla_{\rho}u^{\rho}-x_{2}u^{\rho}\nabla_{\rho}(\beta\mu)$ (55)
$\frac{n_{(1)}}{\beta\hbar}$ $=x_{1}u^{\rho}\nabla_{\rho}\beta-x_{2}\beta\nabla_{\rho}u^{\rho}-y_{1}\beta u^{\rho}\nabla_{\rho}(\beta\mu)$ (56)
$\frac{q_{(1)\sigma}}{\beta\hbar}$ $=w_{3}(\beta^{-1}\nabla_{\sigma}\beta-u^{\rho}\nabla_{\rho}u_{\sigma})-x_{3}(\beta u^{\sigma}F_{\sigma\rho}+\nabla_{\sigma}(\beta\mu)),$ (57)
$\frac{j_{(1)\mu}}{\beta\hbar}$ $=x_{3}(\nabla_{\mu}\beta-\beta u^{\rho}\nabla_{\rho}u_{\mu})-y_{2}(\beta^{2}u^{\nu}F_{\nu\mu}+\beta\nabla_{\mu}(\beta\mu)),$ (58)
Then, the FDRs for the hydrodynamical coefficients are obtained as Crossley:2015evo
$w_{1}$ $=\frac{\epsilon}{\beta\hbar},w_{2}=\frac{\lambda_{1}}{\beta\hbar}=\frac{{\lambda_{2}}}{\beta\hbar},w_{3}=\frac{\kappa_{1}}{\beta\hbar}=\frac{\kappa_{2}}{\beta\hbar},w_{4}=\frac{\zeta}{\beta\hbar},w_{5}=\frac{\eta}{\beta\hbar},$ (59)
$x_{1}$ $=\frac{\nu_{1}}{\beta\hbar}=\frac{\nu_{4}}{\beta\hbar},x_{2}=\frac{\nu_{2}}{\beta\hbar}=\frac{\nu_{5}}{\beta\hbar},x_{3}=\frac{\tau_{1}}{\beta\hbar}=\frac{\tau_{2}}{\beta\hbar}=\frac{\tau_{3}}{\beta\hbar}=\frac{\tau_{4}}{\beta\hbar},$ (60)
$y_{1}$ $=\frac{\nu_{3}}{\beta\hbar},y_{2}=\frac{\chi_{1}}{\beta\hbar}=\frac{\chi_{2}}{\beta\hbar}.$ (61)
For thermal states, we use the following conventional parameters
$\lambda\equiv\lambda_{n},\kappa\equiv\kappa_{n},\tau\equiv\tau_{n},\chi\equiv\chi_{n}.$ (62)
@@ 2.4 Unitarity constraints
The unitarity condition further binds the noise coefficients. Consider 3+1 decomposition of the noise external fields:
$\bar{N}$ $\equiv u^{\mu}u^{\nu}\bar{g}_{\mu\nu},$ (63)
$\bar{N}_{\mu}$ $\equiv u^{\rho}\gamma_{\mu}{}^{\nu}\bar{g}_{\rho\nu},$ (64)
$\bar{\gamma}_{\mu\nu}$ $\equiv\gamma_{\mu}{}^{\rho}\gamma_{\nu}{}^{\sigma}\bar{g}_{\rho\sigma},$ (65)
$\bar{\gamma}$ $\equiv\gamma^{\mu\nu}\bar{\gamma}_{\mu\nu},$ (66)
$\bar{\gamma}_{\perp\mu\nu}$ $\equiv\bar{\gamma}_{\mu\nu}-\frac{\bar{\gamma}}{3}\gamma_{\mu\nu},$ (67)
$\bar{A}$ $\equiv u^{\rho}\bar{A}_{\rho},$ (68)
$\bar{A}_{\perp\mu}$ $\equiv\gamma_{\mu\nu}\bar{A}^{\nu}.$ (69)
@@
$\nu_{1}\bar{A}\bar{N}+\nu_{2}\bar{A}\bar{\gamma}+\nu_{3}\bar{A}^{2}+\chi\bar{A}_{\perp\mu}\bar{A}_{\perp}^{\mu}+2\tau\bar{A}_{\perp\mu}\bar{N}^{\mu}\Bigg{]}.$ (70)
For arbitrary external fields, the second unitarity condition (27) is held. Hence, the following parameter space is allowed Crossley:2015evo :
$\eta$ $\geq 0,$ (71)
$\kappa x^{2}+2\tau xy+\chi y^{2}$ $\geq 0,$ (72)
$\epsilon x^{2}+\zeta y^{2}+\nu_{3}z^{2}+2\lambda xy+2\nu_{1}yz+2\nu_{2}zx$ $\geq 0,$ (73)
with arbitrary real numbers $x,y$, and $z$.
@@ 2.5 Field redefinition
In hydrodynamics, there are ambiguities in the definition of hydrodynamical solutions. Traditionally, such a redefinition of hydrodynamical solutions is called hydrodynamical frame transformation. We only rewrite the same solutions differently, so the generating functional must be unchanged. Hence, genuine physical properties should be independent of such ambiguities. In this section, we summarize the field redefinition dependence of first-order coefficients and find the field redefinition invariant quantities, such as shear viscosity.
Consider redefinition of the hydrodynamical solutions Kovtun:2012rj :
$\beta$ $\to\beta+\delta\beta,$ (74)
$u^{\mu}$ $\to u^{\mu}+\delta u^{\mu},$ (75)
$\mu$ $\to\mu+\delta\mu.$ (76)
$\delta\beta$, $\delta u^{\mu}$ and $\delta\mu$ are generally written as
$\delta\beta$ $=\beta^{-1}\epsilon_{\beta}u^{\rho}\nabla_{\rho}\beta-\lambda_{\beta}\nabla_{\rho}u^{\rho}-\nu_{\beta}u^{\rho}\nabla_{\rho}(\beta\mu),$ (77)
$\delta\mu$ $=\beta^{-1}\epsilon_{\mu}u^{\rho}\nabla_{\rho}\beta-\lambda_{\mu}\nabla_{\rho}u^{\rho}-\nu_{\mu}u^{\rho}\nabla_{\rho}(\beta\mu),$ (78)
$\delta u_{\mu}$ $=\beta^{-1}\kappa_{\beta}\nabla_{\mu}\beta-\kappa_{u}u^{\rho}\nabla_{\rho}u_{\mu}-\tau_{\beta\mu}\nabla_{\mu}(\beta\mu)-\beta\tau_{F}u^{\rho}F_{\rho\mu},$ (79)
with $\delta u^{\mu}=\gamma^{\mu\rho}\delta u_{\rho}$. The redefinition of the hydrodynamical solutions propagates to the hydrodynamical coefficients as follows. First, the zeroth order coefficients are expanded into
$\delta\rho$ $=\rho_{\beta}\delta\beta+\rho_{\mu}\delta\mu,$ (80)
$\delta P$ $=P_{\beta}\delta\beta+P_{\mu}\delta\mu,$ (81)
$\delta n$ $=n_{\beta}\delta\beta+n_{\mu}\delta\mu.$ (82)
Since $\delta\beta$, $\delta u^{\mu}$ and $\delta\mu$ are first order, $\delta\rho$, $\delta P$, and $\delta n$ introduces the transformation for the first order coefficients:
$\delta\epsilon$ $=\rho_{\beta}\epsilon_{\beta}+\rho_{\mu}\epsilon_{\mu},$ (83)
$\delta\lambda_{1}$ $=\rho_{\beta}\lambda_{\beta}+\rho_{\mu}\lambda_{\mu},$ (84)
$\delta\nu_{4}$ $=\rho_{\beta}\nu_{\beta}+\rho_{\mu}\nu_{\mu},$ (85)
$\delta\lambda_{2}$ $=P_{\beta}\epsilon_{\beta}+P_{\mu}\epsilon_{\mu},$ (86)
$\delta\zeta$ $=P_{\beta}\lambda_{\beta}+P_{\mu}\lambda_{\mu},$ (87)
$\delta\nu_{5}$ $=P_{\beta}\nu_{\beta}+P_{\mu}\nu_{\mu},$ (88)
$\delta\nu_{1}$ $=n_{\beta}\epsilon_{\beta}+n_{\mu}\epsilon_{\mu},$ (89)
$\delta\nu_{2}$ $=n_{\beta}\lambda_{\beta}+n_{\mu}\lambda_{\mu},$ (90)
$\delta\nu_{3}$ $=n_{\beta}\nu_{\beta}+n_{\mu}\nu_{\mu}.$ (91)
@@
$\delta q_{\mu}$ $=(\rho+P)\delta u^{\mu},$ (92)
$\delta j_{\mu}$ $=n\delta u^{\mu}.$ (93)
@@
$\delta\kappa_{1}$ $=(\rho+P)\kappa_{\beta},$ (94)
$\delta\kappa_{2}$ $=(\rho+P)\kappa_{u},$ (95)
$\delta\tau_{3}$ $=(\rho+P)\tau_{\beta\mu},$ (96)
$\delta\tau_{4}$ $=(\rho+P)\tau_{F},$ (97)
$\delta\tau_{1}$ $=\frac{n}{\beta}\kappa_{\beta},$ (98)
$\delta\tau_{2}$ $=\frac{n}{\beta}\kappa_{u},$ (99)
$\delta\chi_{1}$ $=\frac{n}{\beta}\tau_{\beta\mu},$ (100)
$\delta\chi_{2}$ $=\frac{n}{\beta}\tau_{F}.$ (101)
The field redefinition does not change the leading order noise coefficients $x_{n}$, $y_{n}$, and $w_{n}$ as $\delta\beta$, $\delta\mu$ and $\delta u^{\mu}$ are first order. Any field redefinition for the KMS symmetric generating functional leads to the violation of the explicit KMS condition with respect to the new $\beta$, which implies that redefined $\beta$ is not the genuine thermodynamical temperature (i.e., the parameter defined in the grand canonical ensemble) at first order. We emphasize that only the explicit KMS condition is not satisfied, and the initial state for the generating functional is unchanged. In fact, the KMS condition is satisfied in a different form Glorioso:2017fpd .
One can reduce the number of hydrodynamical coefficients in compensation for the explicit KMS condition. When maximally eliminating the coefficients with $\rho_{(1)}=n_{(1)}=0$ and $q^{\mu}=0$ the remaining parts are written as
$P_{(1)}$ $=\beta^{-1}\tilde{\lambda}u^{\rho}\nabla_{\rho}\beta-\tilde{\zeta}\nabla_{\rho}u^{\rho}-\tilde{\nu}u^{\rho}\nabla_{\rho}(\beta\mu),$ (102)
$j_{(1)\mu}$ $=\tilde{\tau}_{1}\nabla_{\mu}\beta-\tilde{\tau}_{2}\beta u^{\rho}\nabla_{\rho}u_{\mu}-\tilde{\chi}_{1}\beta\nabla_{\rho}(\beta\mu)-\tilde{\chi}_{2}\beta^{2}u^{\rho}F_{\rho\mu},$ (103)
@@
$\tilde{\lambda}$ $=\lambda_{2}-\left(\frac{\partial P}{\partial\rho}\right)_{n}\epsilon-\left(\frac{\partial P}{\partial n}\right)_{\rho}\nu_{1},$ (104)
$\tilde{\zeta}$ $=\zeta-\left(\frac{\partial P}{\partial\rho}\right)_{n}\lambda_{1}-\left(\frac{\partial P}{\partial n}\right)_{\rho}\nu_{2},$ (105)
$\tilde{\nu}$ $=\nu_{5}-\left(\frac{\partial P}{\partial\rho}\right)_{n}\nu_{4}-\left(\frac{\partial P}{\partial n}\right)_{\rho}\nu_{3},$ (106)
$\tilde{\tau}_{1}$ $=\tau_{1}-\frac{n}{\beta(\rho+P)}\kappa_{1},$ (107)
$\tilde{\tau}_{2}$ $=\tau_{2}-\frac{n}{\beta(\rho+P)}\kappa_{2},$ (108)
$\tilde{\chi}_{1}$ $=\chi_{1}-\frac{n}{\beta(\rho+P)}\tau_{3},$ (109)
$\tilde{\chi}_{2}$ $=\chi_{2}-\frac{n}{\beta(\rho+P)}\tau_{4}.$ (110)
When starting from the most general U(1) charged hydrodynamics with the KMS condition, $\nu_{4}=\nu_{1}$, $\nu_{5}=\nu_{2}$, and Eq. (62) are satisfied (see app. A.2.). In this case, we reduce the number of vector frame invariants to $\tilde{\tau}=\tilde{\tau}_{i}$ and $\tilde{\chi}=\tilde{\chi}_{i}$.

### v3

@@ 2.4 Fluctuation-dissipation relations
The FDRs (2.48) and (2.49) constrain the effective action. The computation of the Lie derivatives of the metric and vector field with respect to $\beta^{\mu}$ is straightforward. We find
$\frac{\rho_{(1)}}{\beta\hbar}$ $=w_{1}\beta^{-1}u^{\rho}\nabla_{\rho}\beta-w_{2}\nabla_{\rho}u^{\rho}-x_{1}u^{\rho}\nabla_{\rho}(\beta\mu),$ (2.51)
$\frac{P_{(1)}}{\beta\hbar}$ $=w_{2}u^{\rho}\beta^{-1}\nabla_{\rho}\beta-w_{4}\nabla_{\rho}u^{\rho}-x_{2}u^{\rho}\nabla_{\rho}(\beta\mu),$ (2.52)
$\frac{n_{(1)}}{\beta\hbar}$ $=x_{1}u^{\rho}\nabla_{\rho}\beta-x_{2}\beta\nabla_{\rho}u^{\rho}-y_{1}\beta u^{\rho}\nabla_{\rho}(\beta\mu),$ (2.53)
$\frac{q_{(1)\sigma}}{\beta\hbar}$ $=w_{3}\left(\beta^{-1}\nabla_{\sigma}\beta-u^{\rho}\nabla_{\rho}u_{\sigma}\right)-x_{3}\left(\beta u^{\sigma}F_{\sigma\rho}+\nabla_{\sigma}(\beta\mu)\right),$ (2.54)
$\frac{j_{(1)\mu}}{\beta\hbar}$ $=x_{3}\left(\nabla_{\mu}\beta-\beta u^{\rho}\nabla_{\rho}u_{\mu}\right)-y_{2}\left(\beta^{2}u^{\nu}F_{\nu\mu}+\beta\nabla_{\mu}(\beta\mu)\right),$ (2.55)
Then, the FDRs for the hydrodynamical coefficients are obtained as [15]
$w_{1}$ $=\frac{\epsilon}{\beta\hbar},\quad w_{2}=\frac{\lambda_{1}}{\beta\hbar}=\frac{\lambda_{2}}{\beta\hbar},\quad w_{3}=\frac{\kappa_{1}}{\beta\hbar}=\frac{\kappa_{2}}{\beta\hbar},w_{4}=\frac{\zeta}{\beta\hbar},\quad w_{5}=\frac{\eta}{\beta\hbar},$ (2.56)
$x_{1}$ $=\frac{\nu_{1}}{\beta\hbar}=\frac{\nu_{4}}{\beta\hbar},\quad x_{2}=\frac{\nu_{2}}{\beta\hbar}=\frac{\nu_{5}}{\beta\hbar},\quad x_{3}=\frac{\tau_{1}}{\beta\hbar}=\frac{\tau_{2}}{\beta\hbar}=\frac{\tau_{3}}{\beta\hbar}=\frac{\tau_{4}}{\beta\hbar},$ (2.57)
$y_{1}$ $=\frac{\nu_{3}}{\beta\hbar},\quad y_{2}=\frac{\chi_{1}}{\beta\hbar}=\frac{\chi_{2}}{\beta\hbar}.$ (2.58)
For thermal states, we use the following conventional parameters:
$\lambda\equiv\lambda_{n},\quad\kappa\equiv\kappa_{n},\quad\tau\equiv\tau_{n},\quad\chi\equiv\chi_{n}.$ (2.59)
@@ 2.5 Unitarity constraints
The unitarity condition further constrains the noise coefficients. Consider the 3+1 decomposition of the noise external fields:
$\bar{N}$ $\equiv u^{\mu}u^{\nu}\bar{g}_{\mu\nu},$ (2.60)
$\bar{N}_{\mu}$ $\equiv u^{\rho}\gamma_{\mu}{}^{\nu}\bar{g}_{\rho\nu},$ (2.61)
$\bar{\gamma}_{\mu\nu}$ $\equiv\gamma_{\mu}{}^{\rho}\gamma_{\nu}{}^{\sigma}\bar{g}_{\rho\sigma},$ (2.62)
$\bar{\gamma}$ $\equiv\gamma^{\mu\nu}\bar{\gamma}_{\mu\nu},$ (2.63)
$\bar{\gamma}_{\perp\mu\nu}$ $\equiv\bar{\gamma}_{\mu\nu}-\frac{\bar{\gamma}}{3}\gamma_{\mu\nu},$ (2.64)
$\bar{A}$ $\equiv u^{\rho}\bar{A}_{\rho},$ (2.65)
$\bar{A}_{\perp\mu}$ $\equiv\gamma_{\mu\nu}\bar{A}^{\nu}.$ (2.66)
@@
$+\nu_{1}\bar{A}\bar{N}+\nu_{2}\bar{A}\bar{\gamma}+\nu_{3}\bar{A}^{2}+\chi\bar{A}_{\perp\mu}\bar{A}_{\perp}^{\mu}+2\tau\bar{A}_{\perp\mu}\bar{N}^{\mu}\Bigg{]}.$ (2.67)
For arbitrary external fields, the second unitarity condition (2.27) is satisfied. Thus, the following parameter space is allowed [15]:
$\eta$ 0$\geq$ 0,≥ 0 , (2.68)
$\kappa x^{2}+2\tau xy+\chi y^{2}$ 0$\geq$ 0,≥ 0 , (2.69)
$\epsilon x^{2}+\zeta y^{2}+\nu_{3}z^{2}+2\lambda xy+2\nu_{1}yz+2\nu_{2}zx$ 0$\geq$ 0,≥ 0 , (2.70)
where $x$, $y$, and $z$ are arbitrary real numbers.
Section summary.
[…]
By imposing the local KMS condition—as a remnant of microscopic time-reversal symmetry and grand canonical ensemble—thermodynamic relations emerge, and the first-order transport coefficients become related to the corresponding noise coefficients. The unitarity condition on the noise sector is then reformulated as constraints on the dissipative coefficients, such as viscosities. In traditional hydrodynamics, these constraints are usually derived from the local second law of thermodynamics imposed phenomenologically; in contrast, our approach derives them directly from the structure of the effective action.
At the outset, the system contains three thermodynamic variables, seventeen first-order dissipative coefficients, and ten zeroth-order noise coefficients. The local KMS condition imposes a thermodynamic relation that allows $\rho$ and $n$ to be expressed in terms of $P$. Its extension to first order—the fluctuation-dissipation relation (FDR)—further relates the seventeen dissipative coefficients to the ten noise coefficients. As a result, the total number of independent parameters is reduced from thirty to eleven  [^2] More rigorously, the equation of state is necessary to fully determine the relationship in thermodynamical variables..
Up to this point, the effective action has been constrained solely by physical principles. Nevertheless, it still includes redundancies associated with field redefinitions—purely unphysical degrees of freedom—which must be properly fixed. These will be addressed in the next section.
3 Frame redundancy and hydrodynamical invariants
When we go to first order theories, there are inherent ambiguities in the definition of $\beta^{\mu}$ at first order. These ambiguities are traditionally referred to as hydrodynamical frame transformations, which correspond to redefinitions of the hydrodynamic fields that leave the physical content unchanged. Since these transformations merely represent different parametrizations of the same underlying dynamics, the generating functional must remain invariant under them. Consequently, physically meaningful results, such as mode analyses, should be expressed in terms of hydrodynamical frame invariants.
In this section, we clarify how the first-order transport coefficients transform under field redefinitions and identify the combinations that remain invariant.
@@ 3.1 Field redefinition
As summarized in the previous section, $\beta^{\mu}$ introduces a rich structure in the constitutive relations. When expanding these relations to first order in derivatives, one may incorporate the expansion parameters into $\beta^{\mu}$ to reduce the number of independent coefficients. This redundancy is associated with $\beta^{\mu}$ itself, rather than with the external fields.
As previously noted, $\mu$ and $\beta$ are projections of the external fields onto $\beta^{\mu}$. Therefore, we cannot redefine both independently. Once we fix $\beta^{\mu}$ by choosing $\delta\beta^{\mu}=u^{\mu}\delta\beta+\beta\delta u^{\mu}$, there is no remaining freedom in $\mu$. Alternatively, we may fix $\delta\beta=0$ and allow $\delta\mu\neq 0$.
As a general choice, one often considers a redefinition of the hydrodynamical fields [1]:
$\beta$ $\to\beta+\delta\beta,$ (3.1)
$u^{\mu}$ $\to u^{\mu}+\delta u^{\mu},$ (3.2)
$\mu$ $\to\mu+\delta\mu.$ (3.3)
$\delta\beta$, $\delta u^{\mu}$, and $\delta\mu$ are generally written as
$\delta\beta$ $=\beta^{-1}\epsilon_{\beta}u^{\rho}\nabla_{\rho}\beta-\lambda_{\beta}\nabla_{\rho}u^{\rho}-\nu_{\beta}u^{\rho}\nabla_{\rho}(\beta\mu),$ (3.4)
$\delta\mu$ $=\beta^{-1}\epsilon_{\mu}u^{\rho}\nabla_{\rho}\beta-\lambda_{\mu}\nabla_{\rho}u^{\rho}-\nu_{\mu}u^{\rho}\nabla_{\rho}(\beta\mu),$ (3.5)
$\delta u_{\mu}$ $=\beta^{-1}\kappa_{\beta}\nabla_{\mu}\beta-\kappa_{u}u^{\rho}\nabla_{\rho}u_{\mu}-\tau_{\beta\mu}\nabla_{\mu}(\beta\mu)-\beta\tau_{F}u^{\rho}F_{\rho\mu},$ (3.6)
with $\delta u^{\mu}=\gamma^{\mu\rho}\delta u_{\rho}$. While ten parameters appear in the expressions above, not all are independent, as discussed earlier.
The redefinition of the hydrodynamical solutions propagates to the hydrodynamical coefficients as follows. First, the zeroth-order coefficients are expanded as
$\delta\rho$ $=\rho_{\beta}\delta\beta+\rho_{\mu}\delta\mu,$ (3.7)
$\delta P$ $=P_{\beta}\delta\beta+P_{\mu}\delta\mu,$ (3.8)
$\delta n$ $=n_{\beta}\delta\beta+n_{\mu}\delta\mu.$ (3.9)
Since $\delta\beta$, $\delta u^{\mu}$, and $\delta\mu$ are first-order terms, $\delta\rho$, $\delta P$, and $\delta n$ introduce the transformation for the first-order coefficients:
$\delta\epsilon$ $=\rho_{\beta}\epsilon_{\beta}+\rho_{\mu}\epsilon_{\mu},$ (3.10)
$\delta\lambda_{1}$ $=\rho_{\beta}\lambda_{\beta}+\rho_{\mu}\lambda_{\mu},$ (3.11)
$\delta\nu_{4}$ $=\rho_{\beta}\nu_{\beta}+\rho_{\mu}\nu_{\mu},$ (3.12)
$\delta\lambda_{2}$ $=P_{\beta}\epsilon_{\beta}+P_{\mu}\epsilon_{\mu},$ (3.13)
$\delta\zeta$ $=P_{\beta}\lambda_{\beta}+P_{\mu}\lambda_{\mu},$ (3.14)
$\delta\nu_{5}$ $=P_{\beta}\nu_{\beta}+P_{\mu}\nu_{\mu},$ (3.15)
$\delta\nu_{1}$ $=n_{\beta}\epsilon_{\beta}+n_{\mu}\epsilon_{\mu},$ (3.16)
$\delta\nu_{2}$ $=n_{\beta}\lambda_{\beta}+n_{\mu}\lambda_{\mu},$ (3.17)
$\delta\nu_{3}$ $=n_{\beta}\nu_{\beta}+n_{\mu}\nu_{\mu}.$ (3.18)
@@
$\delta q_{\mu}$ $=(\rho+P)\delta u^{\mu},$ (3.19)
$\delta j_{\mu}$ $=n\delta u^{\mu}.$ (3.20)
@@
$\delta\kappa_{1}$ $=(\rho+P)\kappa_{\beta},$ (3.21)
$\delta\kappa_{2}$ $=(\rho+P)\kappa_{u},$ (3.22)
$\delta\tau_{3}$ $=(\rho+P)\tau_{\beta\mu},$ (3.23)
$\delta\tau_{4}$ $=(\rho+P)\tau_{F},$ (3.24)
$\delta\tau_{1}$ $=\frac{n}{\beta}\kappa_{\beta},$ (3.25)
$\delta\tau_{2}$ $=\frac{n}{\beta}\kappa_{u},$ (3.26)
$\delta\chi_{1}$ $=\frac{n}{\beta}\tau_{\beta\mu},$ (3.27)
$\delta\chi_{2}$ $=\frac{n}{\beta}\tau_{F}.$ (3.28)
The field redefinition does not change the leading-order noise coefficients $x_{n}$, $y_{n}$, and $w_{n}$ as $\delta\beta$, $\delta\mu$, and $\delta u^{\mu}$ are first-order terms.
Any field redefinition in the KMS-symmetric generating functional generally leads to a violation of the explicit form of the KMS condition with respect to the redefined $\beta^{\mu}$. This implies that the new $\beta^{\mu}$ can no longer be interpreted as the genuine thermodynamic temperature vector that enters the grand canonical ensemble at first order. [^3] We emphasize that only the explicit form of the KMS condition is not preserved; the initial thermal state of the generating functional remains unchanged. The KMS symmetry is still satisfied in a modified, implicit form, as discussed in Ref. [24]. This trade-off allows one to reduce the number of hydrodynamical coefficients at the expense of losing manifest fluctuation-dissipation relations. When maximally eliminating the coefficients with $\rho_{(1)}=n_{(1)}=0$ and $q^{\mu}=0$ the remaining parts are written as
$P_{(1)}$ $=\beta^{-1}\tilde{\lambda}u^{\rho}\nabla_{\rho}\beta-\tilde{\zeta}\nabla_{\rho}u^{\rho}-\tilde{\nu}u^{\rho}\nabla_{\rho}(\beta\mu),$ (3.29)
$j_{(1)\mu}$ $=\tilde{\tau}_{1}\nabla_{\mu}\beta-\tilde{\tau}_{2}\beta u^{\rho}\nabla_{\rho}u_{\mu}-\tilde{\chi}_{1}\beta\nabla_{\rho}(\beta\mu)-\tilde{\chi}_{2}\beta^{2}u^{\rho}F_{\rho\mu},$ (3.30)
@@
$\tilde{\lambda}$ $=\lambda_{2}-\left(\frac{\partial P}{\partial\rho}\right)_{n}\epsilon-\left(\frac{\partial P}{\partial n}\right)_{\rho}\nu_{1},$ (3.31)
$\tilde{\zeta}$ $=\zeta-\left(\frac{\partial P}{\partial\rho}\right)_{n}\lambda_{1}-\left(\frac{\partial P}{\partial n}\right)_{\rho}\nu_{2},$ (3.32)
$\tilde{\nu}$ $=\nu_{5}-\left(\frac{\partial P}{\partial\rho}\right)_{n}\nu_{4}-\left(\frac{\partial P}{\partial n}\right)_{\rho}\nu_{3},$ (3.33)
$\tilde{\tau}_{1}$ $=\tau_{1}-\frac{n}{\beta(\rho+P)}\kappa_{1},$ (3.34)
$\tilde{\tau}_{2}$ $=\tau_{2}-\frac{n}{\beta(\rho+P)}\kappa_{2},$ (3.35)
$\tilde{\chi}_{1}$ $=\chi_{1}-\frac{n}{\beta(\rho+P)}\tau_{3},$ (3.36)
$\tilde{\chi}_{2}$ $=\chi_{2}-\frac{n}{\beta(\rho+P)}\tau_{4}.$ (3.37)
When starting from the most general U(1) charged hydrodynamics with the KMS condition, $\nu_{4}=\nu_{1}$, $\nu_{5}=\nu_{2}$, and Eq. (2.59) are satisfied (see app. B.2.). In this case, we reduce the number of vector frame invariants to $\tilde{\tau}=\tilde{\tau}_{i}$ and $\tilde{\chi}=\tilde{\chi}_{i}$.